6x – 3 = 2x – (4 – 3)/3
Answer by funmath(2933)

Work your parenthesis first:
6x-3=2x-(4-3)/3
6x-3=2x-1/3
Multiply both sides of the equation by 3 to clear your fraction and make life easier.
3(6x-3)=3(2x-1/3)
3(6x)+3(-3)=3(2x)+3(-1/3)
18x-9=6x-3/3
18x-9=6x-1
-6x+18x-9=-6x+6x-1
12x-9=-1
12x-9+9=-1+9
12x=8
12x/12=8/12
x=2/3